OpenStudy (anonymous):

hey struggling and crying my eyes out anyone willing to help me with this assignment

OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = 3(1.02)x x g(x) -1 -4 0 6 1 8 2 10 Which statement best compares the y-intercepts of f(x) and g(x)? (4 points) The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x).

OpenStudy (anonymous):

someone?

OpenStudy (anonymous):

i guess its the second one

OpenStudy (anonymous):

Choose the system of equations which matches the following graph. (4 points) https://invirtual.brainhoney.com/Resource/24528194,A52,0,0/Assets/73608_53bff74f/10_20_q121.jpg x + 2y = −10 3x − y = 12 x − 2y = −10 3x − y = 12 x + 2y = −10 3x + y = 12 x − 2y = −10 3x + y = 12